# MUST-RAG: MUSical Text Question Answering with Retrieval Augmented Generation

### v1

@@ 7 Conclusion
In this paper, we presented MusT-RAG, a retrieval-augmented framework that enhances text-only Music Question Answering (MQA) by adapting general-purpose LLMs to the music domain. By retrieving relevant passages from a music-specific database and incorporating them into the generation context, MusT-RAG effectively mitigates the factuality limitations commonly observed in LLMs. As a result, our method achieves substantial improvements over GPT-4o [35], particularly in factual accuracy. Beyond simple retrieval, we further demonstrated that RAG-style fine-tuning outperforms traditional QA fine-tuning by improving both factual and contextual performance. Our final model achieves a 15.0% gain in factual performance over GPT-4o while maintaining comparable performance in contextual tasks. Importantly, MusT-RAG shows strong generalization capabilities. On the out-of-domain benchmark TrustMus [13], it delivers a 5.7% performance improvement over the zero-shot baseline, underscoring its robustness across diverse music-related QA scenarios. To facilitate future work in this underexplored domain, we release two key resources: MusWikiDB, a music-specific retrieval corpus, and ArtistMus, a benchmark focused on artist-level musical knowledge. We hope these contributions will drive further progress in developing accurate and domain-aware LLMs for music understanding and beyond.

### v2

@@ 7 Conclusion
In this paper, we presented MusT-RAG, a retrieval-augmented framework that enhances text-only Music Question Answering (MQA) by adapting general-purpose LLMs to the music domain. By retrieving relevant passages from a music-specific database and incorporating them into the generation context, MusT-RAG effectively mitigates the factuality limitations commonly observed in LLMs. As a result, our method achieves substantial improvements over GPT-4o [achiam2023gpt], particularly in factual accuracy. Beyond simple retrieval, we further demonstrated that RAG-style fine-tuning outperforms traditional QA fine-tuning by improving both factual and contextual performance. Our final model achieves a 15.0% gain in factual performance over GPT-4o while maintaining comparable performance in contextual tasks. Importantly, MusT-RAG shows strong generalization capabilities. On the out-of-domain benchmark TrustMus [ramoneda2024role], it delivers a 5.7% performance improvement over the zero-shot baseline, underscoring its robustness across diverse music-related QA scenarios. To facilitate future work in this underexplored domain, we release two key resources: MusWikiDB, a music-specific retrieval corpus, and ArtistMus, a benchmark focused on artist-level musical knowledge. We hope these contributions will drive further progress in developing accurate and domain-aware LLMs for music understanding and beyond.